# On the Present Status of Inflationary Cosmology

### v1

@@ 13.2 Pole inflation, D-brane inflation, KKLTI models, polynomial $\alpha$-attractors
$\alpha$-attractors represent a special version of a more general class of attractors, the so-called pole inflation models Galante:2014ifa . It is obtained by generalizing equation (24):
@@
The models with $\beta<2$ provide spectral index $n_{s}$ slightly greater than the $\alpha$-attractors result $n_{s}=1-{2\over N_{e}}$
@@
[…]
@@
$V_{BI}\sim V_{0}(1-{m^{k}\over\varphi^{k}}+...)\ .$ (51)
In Encyclopædia Inflationaris Martin:2013tda , the potentials $V_{BI}\sim V_{0}(1-{m^{k}\over\varphi^{k}})$ were called BI potentials. The numerical evaluation of observables KKLTI models is given in Martin:2013tda for $k=2,3,4$ starting with $r<10^{-1}$. In the subsequent work Kallosh:2018zsi , the stringy origin of the inverse harmonic functions in eq. (46) was clarified, and a detailed comparison of BI and KKLTI models of inflation was presented. It was stressed that at small $r$, predictions of BI and KKLTI models coincide. However, at $r>10^{-3}$ their values of $n_{s}$ may differ significantly. Moreover, it was realized that the values of $n_{s}$ in the KKLTI models have an attractor behavior: their values of $n_{s}$ do not depend on $r$ in the small $r$ limit. The embedding of these models into supergravity was also presented in Kallosh:2018zsi .
Some potentials that appear in the pole inflation scenario may have an alternative interpretation not related to Dp-branes. For example, a quadratic model $V\sim{\varphi^{2}\over m^{2}+\varphi^{2}}$ was first proposed in Stewart:1994pt . It was used in Dong:2010in as an example of a flattening mechanism for the $\varphi^{2}$ potential due to the inflaton interactions with heavy scalar fields. To explain the basic idea, consider a potential
@@
A more general form of the potential which can be used for polynomial $\alpha$-attractors for odd or non-integer $k$ is $V_{0}\,{(\ln^{2}\rho+1)^{k/2}-1\over(\ln^{2}\rho+1)^{k/2}+1}$. In the large $\varphi$ limit, this allows to reproduce the potential (55) for arbitrary positive values of $k$.
@@
Note that for any $k>0$ the attractor values of $n_{s}$ in (56) are greater than $1-2/N_{e}$, and in the limit $k\to 0$ one has $n_{s}\to 1-1/N_{e}$. This means, for example, that for $N_{e}=55$, this class of models can account for $n_{s}$ in the broad range $0.9820.9636<n_{s}<$ 0.982. In Fig. 18, we give a combined plot of the predictions of the simplest exponential $\alpha$-attractor models and polynomial $\alpha$-attractor inflationary models for $n_{s}$ and $\log_{10}r$, for $N_{e}=50$ and $60$ Kallosh:2019hzo . In the small $\mu$ limit, the predicted values of $r$ for all these models can take extremely small values, all the way down to $r\to 0$.
Independently of their interpretation, exponential and polynomial $\alpha$-attractors may serve as a powerful tool for parametrization of all observational data with $1-2/N_{e}\lesssim n_{s}<1-1/N_{e}$, i.e. in the range $0.980.96<n_{s}<$ 0.98 for $N_{e}=50$, or $0.967\lesssim n_{s}\lesssim 0.983$ for $N_{e}=60$ Kallosh:2019eeu ; Kallosh:2019hzo .
@@
Figure 19: An extended version of the figure from “Snowmass2021 Cosmic Frontier: Cosmic Microwave Background Measurements White Paper” Chang:2022tzj which was presented in Fig. 1 in Kallosh:2022vha . It shows the predictions of T-model $\alpha$-attractors with unconstrained values of $\alpha$ (gray area), E-models (red lines), the predictions for $13\alpha=7,6,5,4,3,2,$ 1 (purple lines), as well as Higgs inflation, $R^{2}$ inflation (red dots). In addition, the purple lines show the prediction of the polynomial attractors (47), and the orange lines correspond to (48). The predictions are for $5747<N_{e}<$ 57.
Finally, we should mention a distinguishing feature of the polynomial $\alpha$-attractors. Even though these models describe the potentials of the same type (55) as all other models described in this section, the polynomial $\alpha$-attractor have the same geometric origin as all other $\alpha$-attractor models, and therefore they also have a discrete series of preferred values $73\alpha=1,2,...,$ 7. As a result, for each value of $k$, we have a series of predictions for $r$ corresponding to these values of $3\alpha$. For example, in the simplest model with $k=2$ we have the following Poincaré disk series:
@@ 14 Hybrid attractors
In the original version of this model, uplift is large, $n_{s}$ typically is very close to 1, and $r$ can be extremely small. Indeed, expressions for $1-n_{s}$ and $r$ contain terms proportional to $(V^{\prime}/V)^{2}$ and $V^{\prime\prime}/V$. For any given $\phi$, the uplift $V_{\rm up}$ does not change the derivatives of the potential while increasing $V$, which decreases $1-n_{s}$ and $r$. But if the uplift is very small, inflation occurs as in the simple chaotic inflation model ${m^{2}\over 2}\phi^{2}$, and $n_{s}=1-2/N_{e}$.
A similar situation appears if one embeds the model (59) in the context of $\alpha$-attractors Kallosh:2022ggf . In this class of models, one has two basic attractor regimes. If $V_{\rm up}$ is large and $\phi_{c}\ll 1$, then the last stage of inflation occurs as in the original hybrid inflation scenario, with $n_{s}$ very close to 1. On the other hand, if uplifting is small, one has the standard $\alpha$-attractor result $n_{s}=1-{2\over N_{e}},r={12\over N^{2}_{e}}$. Thus, by changing the parameters of the model, one may interpolate between $n_{s}=1-{2\over N_{e}}$ and the $n_{s}=1$ Kallosh:2022ggf .
The full theory of this effect is rather nuanced as it depends on relations between many parameters of the $\alpha$-attractor version of the model (59). For example, in certain cases, the process of spontaneous symmetry breaking can drive a short second stage of inflation, resulting in the production of large primordial black holes Braglia:2022phb , which might even account for dark matter Garcia-Bellido:1996mdl . For some other relations between the parameters, the minimum of the potential at $\sigma=M/\sqrt{\lambda}$ can never be reached, and the standard Higgs type symmetry breaking potential of the field $\sigma$ transforms into a quintessential inflation potential shown in the right part of Fig. 4 in Kallosh:2022ggf .
@@ 15 String theory inflation
Here we will make a few comments, in particular, with regard to the status of more recent string theory inflation models and recent observational results.
@@ 1. Fibre inflation, see Cicoli:2024bxw and references therein, has the effective potential
Even if confirmed by future observations, this model will not unambiguously support string theory “from the sky”, being degenerate with $\alpha$-attractors for $\alpha=2$. Moreover, in both cases, there is an issue raised by the ACT constraint $n_{s}=0.9743\pm 0.0034$. Future data on $n_{s}$ from SPT-3G and from Simons Observatory, as well as theoretical developments, will be very important in this respect.
@@ 16 Discussion
In this short review, we have mentioned many different inflationary models, and yet there are many others which can be found in the Encyclopædia Inflationaris Martin:2013tda . It is difficult to make any definite conclusions based on our subjective choice of the models to discuss, and it is especially difficult in a situation where there is an ongoing discussion of the $H_{0}$ problem and of the validity of the basic $\Lambda$CDM model.
But we will try to do our best, and we will update this article if the new data forces us to reconsider our conclusions. We will divide this Discussion into several different parts.
@@
However, one of our goals pursued in this paper was to describe simple but well-motivated models that could describe all CMB-related data by a choice of a single parameter. Also, the polynomial model mentioned above is a single-field model, as are most other models discussed in this paper. A discovery of local non-Gaussianity $f_{NL}>O(1)$ would force us to consider more complicated models. This explains the importance of investigating non-Gaussianity, see e.g. Heinrich:2023qaa ; Shiveshwarkar:2023afl .
@@
The string inflation models have been tested already. As we mentioned in Sec. 15, 6 models out of the 14 models presented in the table 20 are already ruled out. As for the remaining models, one can identify two different possibilities if inflationary tensor modes are discovered.
@@
b) Suppose instead that observations find tensor modes with any value of $r$ in the very broad range $0.03610^{{-4}}<r<$ 0.036, but far from the point $n_{s}=0.963$, $r\approx 0.0072$ favored by Fibre inflation. Rather paradoxically, this discovery would rule out all string inflation models mentioned in the review article Cicoli:2023opf and in our paper. This “boom-or-bust” scenario is yet another reason to continue searching for tensor modes.
@@
Most of the models that attempt to match the recent ACT constraint $n_{s}=0.9743\pm 0.0034$ have plateau potentials where the potential approaches the plateau not exponentially but polynomially. For example, for polynomial $\alpha$-attractors Kallosh:2022feu
@@
By increasing the precision of determination of $n_{s}$ one may be able to distinguish the two different classes of inflationary models with qualitatively different families of potentials (66) and (67). Note that Simons Observatory plans to halve the current error bar on the scalar perturbation spectral index $n_{s}$ SimonsObservatory:2025wwn .
Precise determination of $n_{s}$ could be sufficient for distinguishing between different classes of models even if they predict too small values of $r$ to be ever observed, because the predictions of $n_{s}$ beciome increasingly accurate in the small $r$ limit, see Fig. 18. This may help us to distinguish not only between the general classes of models (66) and (67), but also between different ${\rm Dp}-{\overline{\rm Dp}}$ brane inflation plateau potentials (46) and the Loop Blow-up Inflation (63).
@@
As we already mentioned, there are many other models that deserve to be discussed in this review; see Encyclopædia Inflationaris for a much longer and yet unavoidably incomplete list, each of the models is there for a good reason Martin:2013tda . Moreover, the discussion so far only briefly mentioned potential seismic shifts that are still possible in cosmology. The examples are abundant: the $H_{0}$ problem, dark energy models challenging $\Lambda$CDM, concerns about possible anisotropies of the cosmological background, etc. In this ever-changing situation, it is encouraging that some inflationary models developed long ago still continue working.
@@ Acknowledgement
We acknowledge stimulating discussions with R. Bond, C. Burness, J.J. Carrasco, M.Cortês, G. Efstathiou, F. Finelli, W. Freedman, C. Hill, C.L. Kuo, A. Liddle, V. Mukhanov, F. Quevedo, D. Roest, E. Silverstein, T. Wrase, and Y. Yamada. This work was supported by SITP and by NSF Grant PHY-2310429. It is an extended version of the talk given at the Lemaître Conference 2024 “Black Holes, Gravitational Waves and Space-Time Singularities” at Specola Vaticana. We are grateful to the organizers for their support and hospitality.

### v3

@@ 13.2 Pole inflation, D-brane inflation, KKLTI models, polynomial $\alpha$-attractors
$\alpha$-attractors represent a special version of a more general class of attractors, known as pole inflation models Galante:2014ifa . It is obtained by generalizing equation (26):
@@
The models with $\beta<2$ provide spectral index $n_{s}$ slightly greater than the $\alpha$-attractors result $n_{s}=1-{2\over N_{e}}$ .
@@
Their attractor formula for the $n_{s}$ is given in (47), whereas the formula for $r$ depends on the parameter $m$ in the potential. For ${\rm D3}-{\overline{\rm D3}}$ inflation for small $m$ one has
@@
[…]
In Encyclopædia Inflationaris Martin:2013tda , the potentials $V_{BI}\sim V_{0}(1-{m^{k}\over\varphi^{k}})$ were called BI potentials.
The numerical evaluation of observables KKLTI models is given in Martin:2013tda for $k=2,3,4$ starting with $r<10^{-1}$. In the subsequent work Kallosh:2018zsi , the stringy origin of the inverse harmonic functions in eq. (48) was clarified, and a detailed comparison of BI and KKLTI models of inflation was presented. It was emphasized that at small $r$, the predictions of the BI and KKLTI models coincide. However, at $r>10^{-3}$ their values of $n_{s}$ may differ significantly. Moreover, it was realized that the values of $n_{s}$ in the KKLTI models have an attractor behavior: their values of $n_{s}$ do not depend on $r$ in the small $r$ limit. The embedding of these models into supergravity was also presented in Kallosh:2018zsi , see also Pallis:2025gii .
Some potentials that appear in the pole inflation scenario may have an alternative interpretation unrelated to Dp-branes. For example, a quadratic model $V\sim{\varphi^{2}\over m^{2}+\varphi^{2}}$ was first proposed in Stewart:1994pt . It was used in Dong:2010in as an example of a flattening mechanism for the $\varphi^{2}$ potential due to the inflaton interactions with heavy scalar fields. To explain the basic idea, consider a potential
@@
A more general form of the potential which can be used for polynomial $\alpha$-attractors for odd or non-integer $k$ is $V_{0}\,{(\ln^{2}\rho+1)^{k/2}-1\over(\ln^{2}\rho+1)^{k/2}+1}$. In the large $\varphi$ limit, this allows us to reproduce the potential (57) for arbitrary positive values of $k$.
@@
Note that for any $k>0$ the attractor values of $n_{s}$ in (58) are greater than $1-2/N_{e}$, and in the limit $k\to 0$ one has $n_{s}\to 1-1/N_{e}$. This means, for example, that for $N_{e}=55$, this class of models can account for $n_{s}$ in the broad range $0.9636<n_{s}<0.982$. In Fig. 18, we give a combined plot of the predictions of the simplest exponential $\alpha$-attractor models and polynomial $\alpha$-attractor inflationary models for $n_{s}$ and $\log_{10}r$, for $N_{e}=50$ and $60$ Kallosh:2019hzo . In the small $\mu$ limit, the predicted values of $r$ for all these models can take extremely small values, all the way down to $r\to 0$.
Independently of their interpretation, exponential and polynomial $\alpha$-attractors may serve as a powerful tool for parametrization of all observational data with $1-2/N_{e}\lesssim n_{s}<1-1/N_{e}$, i.e. in the range $0.96<n_{s}<0.98$ for $N_{e}=50$, or $0.967\lesssim n_{s}\lesssim 0.983$ for $N_{e}=60$ Kallosh:2019eeu ; Kallosh:2019hzo .
@@
Figure 19: An extended version of the figure from “Snowmass2021 Cosmic Frontier: Cosmic Microwave Background Measurements White Paper” Chang:2022tzj which was presented in Fig. 1 in Kallosh:2022vha . It shows the predictions of T-model $\alpha$-attractors with unconstrained values of $\alpha$ (gray area), E-models (red lines), the predictions for $3\alpha=7,6,5,4,3,2,1$ (purple lines), as well as Higgs inflation, $R^{2}$ inflation (red dots). In addition, the purple lines show the prediction of the polynomial attractors (49), and the orange lines correspond to (50). The predictions are for $47<N_{e}<57$.
Finally, we should mention a distinguishing feature of the polynomial $\alpha$-attractors. Even though these models describe the potentials of the same type (57) as all other models described in this section, the polynomial $\alpha$-attractor have the same geometric origin as all other $\alpha$-attractor models, and therefore they also have a discrete series of preferred values $3\alpha=1,2,...,7$. As a result, for each value of $k$, we have a series of predictions for $r$ corresponding to these values of $3\alpha$. For example, in the simplest model with $k=2$ we have the following Poincaré disk series:
@@ 14 Hybrid attractors
In the original version of this model, the uplift is large, $n_{s}$ is typically very close to 1, and $r$ can be extremely small. Indeed, expressions for $1-n_{s}$ and $r$ contain terms proportional to $(V^{\prime}/V)^{2}$ and $V^{\prime\prime}/V$. For any given $\phi$, the uplift $V_{\rm up}$ does not change the derivatives of the potential while increasing $V$, which decreases $1-n_{s}$ and $r$. But if the uplift is very small, inflation occurs as in the simple chaotic inflation model ${m^{2}\over 2}\phi^{2}$, and $n_{s}=1-2/N_{e}$.
A similar situation appears if one embeds the model (61) in the context of $\alpha$-attractors Kallosh:2022ggf . In this class of models, one has two basic attractor regimes. If $V_{\rm up}$ is large and $\phi_{c}\ll 1$, then the last stage of inflation occurs as in the original hybrid inflation scenario, with $n_{s}$ very close to 1. On the other hand, if uplifting is small, one has the standard $\alpha$-attractor result $n_{s}=1-{2\over N_{e}},r={12\over N^{2}_{e}}$. Thus, by changing the parameters of the model, one may interpolate between $n_{s}=1-{2\over N_{e}}$ and the $n_{s}=1$ Kallosh:2022ggf ; Iacconi:2024hmg .
The full theory of this effect is rather nuanced as it depends on relations between many parameters of the $\alpha$-attractor version of the model (61). For example, in certain cases, the process of spontaneous symmetry breaking can drive a short second stage of inflation, resulting in the production of large primordial black holes Braglia:2022phb , which might even account for dark matter Garcia-Bellido:1996mdl . For some other relations between the parameters, the minimum of the potential at $\sigma=M/\sqrt{\lambda}$ can never be reached, and the standard Higgs-type symmetry-breaking potential of the field $\sigma$ transforms into a quintessential inflation potential shown in the right part of Fig. 4 in Kallosh:2022ggf .
@@ 15 String theory inflation
Here, we will make a few comments regarding the status of recent string theory inflation models and observational results.
@@ 1. Fibre inflation, see Cicoli:2024bxw and references therein, has the effective potential
Even if confirmed by future observations, this model will not unambiguously support string theory “from the sky”, being degenerate with $\alpha$-attractors for $\alpha=2$. Moreover, in both cases, there is an issue raised by the ACT constraint $n_{s}=0.9743\pm 0.0034$. Future data on $n_{s}$ from SPT-3G and the Simons Observatory, as well as theoretical developments, will be crucial in this respect.
@@ 16 Discussion
[…]
However, we will do our best and update this article if new data forces us to reconsider our conclusions. We will divide this Discussion into several different parts.
@@
However, one of our goals pursued in this paper was to describe simple but well-motivated models that could describe all CMB-related data by a choice of a single parameter. Also, the polynomial model mentioned above is a single-field model, as are most other models discussed in this paper. A discovery of local non-Gaussianity $f_{NL}>O(1)$ would force us to consider more complicated models. This highlights the importance of investigating non-Gaussianity; see, e.g., Heinrich:2023qaa ; Shiveshwarkar:2023afl .
@@
The string inflation models have already been tested. As we mentioned in Sec. 15, 6 models out of the 14 models presented in the table 20 are already ruled out. Regarding the remaining models, two distinct possibilities emerge if inflationary tensor modes are discovered.
@@
b) Suppose instead that observations find tensor modes with any value of $r$ in the very broad range $10^{{-4}}<r<0.036$, but far from the point $n_{s}=0.963$, $r\approx 0.0072$ favored by Fibre inflation. Rather paradoxically, this discovery would rule out all string inflation models mentioned in the review article Cicoli:2023opf .
This does not mean that it is impossible to construct a string theory model consistent with the ACT constraint and predicting $r$ in the range $10^{{-4}}<r<0.036$. An example of such a single-inflaton model was proposed recently in Leontaris:2025hly . This model relies on a judicious choice of more than ten different parameters.
@@
Many models attempting to match the recent ACT constraint $n_{s}=0.9743\pm 0.0034$ have plateau potentials where the potential approaches the plateau not exponentially, but rather polynomially. For example, for polynomial $\alpha$-attractors Kallosh:2022feu
@@
By increasing the precision of determination of $n_{s}$, one may be able to distinguish the two different classes of inflationary models with qualitatively different families of potentials (68) and (69). Note that Simons Observatory plans to halve the current error bar on the scalar perturbation spectral index $n_{s}$ SimonsObservatory:2025wwn .
Precise determination of $n_{s}$ could be sufficient for distinguishing between different classes of models even if they predict too small values of $r$ to be ever observed, because the predictions of $n_{s}$ become increasingly accurate in the small $r$ limit, see Fig. 18. This may help us to distinguish not only between the general classes of models (68) and (69), but also between different ${\rm Dp}-{\overline{\rm Dp}}$ brane inflation plateau potentials (48) and the Loop Blow-up Inflation (65).
@@
[…]
@@ Acknowledgement
We acknowledge stimulating discussions with R. Bond, C. Burness, J.J. Carrasco, M.Cortês, G. Efstathiou, F. Finelli, W. Freedman, C. Hill, C.L. Kuo, A. Liddle, V. Mukhanov, F. Quevedo, D. Roest, E. Silverstein, T. Wrase, and Y. Yamada. This work was supported by the Leinweber Institute for Theoretical Physics at Stanford and by NSF Grant PHY-2310429. It is an extended version of the talk given at the Lemaître Conference 2024 “Black Holes, Gravitational Waves and Space-Time Singularities” at Specola Vaticana. We are grateful to the organizers for their support and hospitality.
Appendix A General multi-field inflationary equations for $\phi^{i}(N)$
Most of inflationary models we discuss in this paper are single field inflationary models. However, they often originate from multi-field models with curved space geometry. We have found it useful to describe background inflationary trajectories for such models either as evolution in time, $\phi^{i}(t)$ or as evolution in $N$, $\phi^{i}(N)$. Here the number of e-foldings is defined as $N\equiv\ln{a(t_{\rm end})\over a(t)}$.
The standard inflationary background equations in theories with an arbitrary curved scalar field geometry and a space-time FRW metric are well known. An advantage of using these equations with evolution in $N$ shows up, in particular, in numerical studies of single field models where potentials are simple for non-canonical fields. Examples in this paper where we used this setup are in Secs. 8, 13.1.
Consider cosmological models in FRW space-time where the field space has a general type curved geometry ${g}_{ij}(\phi)$
$\sqrt{-g}^{-1}\mathcal{L}=\left[-\frac{1}{2}R+\frac{1}{2}{g}_{ij}(\phi)\,\partial_{\mu}\phi^{i}\partial^{\mu}\phi^{j}-V(\phi)\right]\;.$ (70)
Exact background equations of motion for the homogeneous fields $\phi^{i}(t)$ are
$\ddot{\phi}^{i}+3H\dot{\phi}+\Gamma^{i}_{jk}\dot{\phi}^{j}\dot{\phi}^{k}+g^{ij}{\partial V\over\partial\phi^{j}}=0\ ,$ (71)
where
$H(t)\equiv{\dot{a}\over a}=\sqrt{V+E_{\rm kin}\over 3}\,,\qquad E_{\rm kin}=\frac{1}{2}{g}_{ij}(\phi)\,\partial_{\mu}\phi^{i}\partial^{\mu}\phi^{j}\,.$ (72)
These equations were presented in Kallosh:2002gf (we skip here the cold dark matter energy density).
The standard form of inflationary slow-roll parameters can be given in the form Lyth:2009imm ; Iacconi:2021ltm
$\epsilon_{H}=-{\dot{H}\over H^{2}}\,,\qquad\eta_{H}=\epsilon_{H}-{1\over 2}{\dot{\epsilon}_{H}\over H\epsilon_{H}}\ .$ (73)
For single-field models, the observables are
$n_{s}=1-4\epsilon_{H}+2\eta_{H}\ ,\qquad r=16\epsilon_{H}\ .$ (74)
To switch from time evolution to the number of e-foldings $N$ evolution, we use a relation
$dN\equiv d\ln a=Hdt\ ,$ (75)
We use for $\partial\over\partial N$} derivatives, instead of dot for $\partial\over\partial t$}. The inflationary slow-roll parameters are
$\epsilon_{H}=-{H^{\prime}\over H}\,,\qquad\eta_{H}=\epsilon_{H}-{1\over 2}{\epsilon^{\prime}_{H}\over\epsilon_{H}}\ ,$ (76)
and observables depend on the number of e-foldings $N$. After a change of variables according to Eq. (75) our inflationary background equations with $N$ as an evolution parameter take the form
$\phi^{\prime\prime}{}^{i}+(3-\epsilon_{H})\phi^{\prime i}+\Gamma^{i}_{jk}\phi^{\prime j}\phi^{\prime k}+{g^{ij}\over H^{2}}{\partial V\over\partial\phi^{j}}=0\ .$ (77)
By solving it one can find a solution $\phi^{i}(N)$. These equations are supplemented by the following relations
$H^{2}={V\over 3-\epsilon_{H}}\,,\qquad\epsilon_{H}=-{H^{\prime}\over H}=\frac{1}{2}{g}_{ij}(\phi)\,\phi^{\prime i}\phi^{\prime j}\ .$ (78)
Analogous equations were presented in Peterson:2010np where it was also shown that in the case of the two-field models, the extraction of observables like $n_{s},r$ is complicated and depends on the turn rate of the 2-field background trajectory.
Simple examples in single field case:
For the case of models with a nonminimal coupling to gravity Kallosh:2013tua described here in Sec. 8, one can use eq. (15), where the single scalar $\phi$ in the Einstein frame is not canonical, but the potential is given by a well-defined function of the original field $\phi$. Namely, the kinetic term of the scalars involves a geometry $g_{11}d\phi^{1}d\phi^{1}$
$V_{E}={V_{J}\over\Omega^{2}}\ ,\qquad g_{11}=\Omega^{-1}+{3\over 2}(\log\Omega)^{{}^{\prime}2}\ .$ (79)
Thus, we have a single evolution equation where $\phi$ depends on $N$:
$\phi^{\prime\prime}+(3-\epsilon_{H})\phi^{\prime}+\Gamma^{1}_{11}\phi^{\prime}\phi^{\prime}+{g^{11}\over H^{2}}{\partial V\over\partial\phi}=0\,,\qquad\Gamma^{1}_{11}={1\over 2}\partial_{\phi}(\ln(g_{11}))\ .$ (80)
This case was studied for models with $\Omega=1+\xi\phi^{n/2}$ in Kallosh:2013tua . It is very simple to calculate $n_{s}$ and $r$ using the method described above. This procedure was also used for the model of the chaotic inflation with nonminimal coupling to gravity in Sec. (40).